jack mows the lawn in 7 hours and susie can mow the same lawn in 4 hours, how long would it take them to mow that lawn together?
Found 2 solutions by Alan3354, RAY100:
Answer by Alan3354(69443)

jack mows the lawn in 7 hours and susie can mow the same lawn in 4 hours, how long would it take them to mow that lawn together?
------------------
You have to get it into lawns per hour, then add them up.
Jack does 1/7 per hour, Susie 1/4.
1/4 + 1/7 = 11/28 lawns per hour.
That makes 28/11 hours per lawn.
--------------------
There's a shortcut: product over sum
7*4/(7+4) = 28/11

Answer by RAY100(1637)

let x = total job,,, sq ft etc
jacks rate is ,,,,x/7,,,,Sues rate is x/4
rate *time = job
x/7*t+x/4*t=x
mult by 28,,,,divide by x
4t+7t=28
11t=28
t=28/11=2.55 hrs ANSWER
check
x/7 *2.55+x/4*2.55=x
.364x +.636x =x
x=x ok